Hey please help me out in this personal business related question!(wholesale)

Take in consideration prior to this question I have done benchmarking in prices.


Ok so im selling this product that happily costs me to produce about $5 each piece and I’m selling it for $8 each piece to stores in which I do free delivery too since I deliver only here in the state area. I make delivery amounts of $350, $500, and even $850 daily with some of them paying net30 and others instantly...
But is very stressful to make this as I also pay my only 3 workers individual $400 a week for 10 hrs each day but hence they do indeed produce me about 400 pieces a day.

My question is;
1)Am i selling it to cheap?
2)Should I Eliminate a factor : Free deliver?
Net 30 days?
To much employers?
3)Never make them by a amount of $500 or less?


Because customers want more and more as they keep calling me but then again I don’t know if I’m making enough profit for my $2000 a month expenses as I feel I should be making more money even though Is carrying me for a living as of now.
Answer #1

1. Here is my working to determine whether you are selling it too cheap.

What I have understood is that your revenue per day sometime is $350, sometimes it is $500, and sometimes it is $850.

On average let's take it to be $500 a day it means your contribution or we can say saving after deducting cost to produce for 1 day will be around (500/8)x(8-5) = $187.5

So weekly saving before paying to worker = 187.5x7 = $1312.5

Net earning = $1312.5-(400x3) = $112.5 per week which is just a meager amount. In that too delivery expenses like petrol cost etc. is also be deducted.

Now if you price it a bit higher say $9, then the situation would be somehow would be

Total delivery amount = (500/8)x9 = 562.5 i.e. more by $62.5

Therefore profit in that case would be $175 per week, which would be better. Further increase can lead to loss of many customers . Which will not be advisable .

Therefore , yes you are selling it cheap currently and can increase the price to $9 which will not hurt the customers. In fact you can offer them if they buy in bulk (delivery amount you can decide), then you can also give them discount of upto 4% which is called volume discount. It will increase your sale and will lead to more profitability.

2 . I think you should not eliminate these factor it is because these offers looks very lucrative and attracts customer . Rather you can have some condition relating to distance like upto that much kilometers there will be no charges. Net 30 days is fine there is no problem in it infsct for prompt payment you can offer them some discount.

3 . Do not do this because if currently someone is buying below $500 , may in future start buying even more than $500. But if you do not allow them and they shift to some other seller then you may start losing customers.